Exercise 2-15 Plantwide and Departmental Predetermined Overhead Rates; Job Costs [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 30,000 40,000 70,000
Fixed manufacturing overhead costs $ 760,000 $ 210,000 $ 970,000
Variable manufacturing overhead cost per machine-hour $ 5.30 $ 5.30

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 376,000 $ 330,000 $ 706,000
Direct labor cost $ 210,000 $ 160,000 $ 370,000
Machine-hours 24,000 6,000 30,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 250,000 $ 260,000 $ 510,000
Direct labor cost $ 180,000 $ 250,000 $ 430,000
Machine-hours 6,000 34,000 40,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

Exercise 2-15 Part 1

Required:

1. Assume Delph uses a plantwide predetermined overhead rate based on machine-hours.

a. Compute the plantwide predetermined overhead rate.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 150% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?
Answer #1

a. Fixed manufacturing pre determined overhead rate = Fixed manufacturing overhead costs / Machine-hours

= 970,000 / 70,000

= 13.85

plantwide predetermined overhead rate = 13.85+ 5.3 = 19.16 ( rounded to two decimal places)

b.

D70 C 200
Direct materials cost 706000 510000
Direct labor cost 370000 430000
Manufacturing overhead 574800 766400
[30000*19.16] [40000*19.16]
Total manufacturing costs 1650800 1706400

c.

D70 C 200
Total manufacturing costs 1650800 1706400
Mark up (150%) 2476200 2559600
Bid Price 4127000 4266000

d.

Cost of goods sold = 1650800+1706400 = 3357200